OpenStudy (anonymous):

what is the derivative of y=7csc4x?

OpenStudy (johnweldon1993):

use the chain rule...you know you can take 7 out of the equation since It is a constant multiple...so firstly...what is the derivative of csc?

OpenStudy (anonymous):

would it be cscxcotx?

OpenStudy (johnweldon1993):

it would be -csc(x)cot(x) right? don't forget the '-' but yes.... now THAT is the derivative of csc(x)...but what YOU have...is csc(4x) so that cannot be correct...yet so far you have 7 * -csc(4x)(cot(4x) * d/dx 4x which you know is 4 so altogether you have 7 * -csc(4x)cot(4x) * 4 which comes out to 28 (-csc(4x)(cot(4x)) or -28(csc(4x)(cot(4x))

OpenStudy (anonymous):

oh yes, i had forgotten the '-'

OpenStudy (johnweldon1993):

And there you go :) you see what I did correct?

OpenStudy (anonymous):

yes, thanks

OpenStudy (johnweldon1993):

Also ...an EASY way to remember which derivatives has a '-' in them d/dx cos(x) = -sin(x) d/dx cot(x) = -csc²(x) d/dx csc(x) = -cot(x)csc(x) see how each terms that starts with a 'c' has a NEGATIVE derivative? Cos Cot = all negative derivatives :) Csc

OpenStudy (anonymous):

oh yea, thanks again

OpenStudy (johnweldon1993):

anytime!